OpenStudy (anonymous):

The ratio of men to women in a certain factory is 3 to 4. There are 210 men. How many workers are there? @hartnn @taylee1 @TTP @misz.lopez115

hartnn (hartnn):

so lets say there are x total workers men + women = x men/ women = 3/ 4 men = 210 so 210/ women = 3/4 can you first find number of women from here ?

OpenStudy (anonymous):

the quick way to do this sort of problem is to take 210, divide by 3, and then multiply by 4. That will give you the number of women. Add 210 to that number and you have the total number of workers.

OpenStudy (the_fizicx99):

3/4 = 210/x --> 210 * 3 / 4

OpenStudy (anonymous):

So 280 women @misz.lopez115 and @hartnn

hartnn (hartnn):

YES

hartnn (hartnn):

so total = 210 +280 = ..

OpenStudy (anonymous):

yes :) good job!

OpenStudy (anonymous):

Yup, thanks @misz.lopez115 and @hartnn :) big help!!!

OpenStudy (anonymous):

welcome <3